OpenStudy (anonymous):

Solve and show all steps:

OpenStudy (anonymous):

\[y - \frac{ 1 }{ 3 } = \frac{ 2 }{ 9 }\]

OpenStudy (anonymous):

y-1/3=2/9 y=2/9+1/3 or 2/9 + 3/9 y=5/9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem, Glad I can help!